OpenStudy (anonymous):

Part 1: Find the solution to the equation 64(2 – x) = 44x. (4 points) Part 2: Using complete sentences, explain the procedure used to solve this equation. (4 points)

OpenStudy (anonymous):

x=128/108

OpenStudy (asnaseer):

\[64(2-x)=44x\]first simplify by dividing both sides by their common factor (which is 4) \[16(2-x)=11x\]then expand the bracketed expression on the left \[32-16x=11x\]then you need to place all terms with x on one side and all constants on the other side. so we simply move the -16x from the left-hand-side to the right-hand-side thereby changing its sign \[32=11x+16x=27x\]now divide both sides by 27 to get an expression for x \[\frac{32}{27}=x\]re-write in standard form with the subject on the left \[x=\frac{32}{27}\]

OpenStudy (anonymous):

128-64x=44x, 128=108x, x=128/108, x=32/27

OpenStudy (anonymous):

Thank you